# CulturalBench: a Robust, Diverse and Challenging Benchmark on Measuring the

### v1

@@ Acknowledgments
Appendix A Results
Appendix B CulturalBench statistics
Country Counts
North America ($N=27$)
Canada 7
United States 20
South America ($N=150$)
Argentina 35
Brazil 25
Chile 22
Mexico 49
Peru 19
East Europe ($N=115$)
Czech Republic 25
Poland 24
Romania 15
Ukraine 21
Russia 30
South Europe ($N=76$)
Spain 40
Italy 36
West Europe ($N=96$)
France 14
Germany 32
Netherlands 11
United Kingdom 25
Africa ($N=134$)
Egypt 20
Morocco 17
Nigeria 22
South Africa 58
Zimbabwe 17
Country Counts
Middle East/West Asia ($N=127$)
Iran 37
Israel 13
Lebanon 22
Saudi Arabia 17
Turkey 38
South Asia ($N=106$)
Bangladesh 25
India 46
Nepal 21
Pakistan 14
Southeast Asia ($N=159$)
Indonesia 26
Malaysia 11
Philippines 45
Singapore 23
Thailand 27
Vietnam 27
East Asia ($N=211$)
China 59
Hong Kong 36
Japan 53
South Korea 41
Taiwan 22
Oceania ($N=26$)
Australia 15
New Zealand 11
Table 2: Country distribution of 45 countries in CulturalBench
Figure 8: 45 Countries covered by CulturalBench and 16 countries covered by Blend.
Appendix C Definition and examples of cultural topics
@@
Table 3: Examples of CulturalBench for each topic.
@@
Table 4: (conc.) Examples of CulturalBench for each topic.
@@ Appendix D Zero-shot evaluation prompts
Our evaluation is to ask the model in multiple choice setting. The zero-shot prompt is as follow to ensure the model only output one label (A, B, C or D).
@@
Our evaluation is to ask the model in binary setting (True/False). Our prompt is as follow to ensure the model only output one label (True/False).
@@ Appendix E Evaluation results on CulturalBench-Easy
Models North America South America East Europe South Europe West Europe Africa Middle East/ West Asia South Asia Southeast Asia East Asia Oceania
gpt-3-5-turbo-1106 80.72 71.34 72.64 83.75 78.61 81.22 54.71 80.68 67.5 63.7 86.36
gpt35turbo 87.85 72.21 76.86 80.97 74.61 80.74 59.19 83.82 68.39 62.04 78.48
gpt4omini 90.35 78.54 89.08 89.16 87.31 87.24 71.4 90.46 81.2 79.14 87.57
gpt4o 100.0 87.82 91.93 91.66 97.22 91.81 80.78 90.13 86.07 86.44 92.12
gpt-4o-2024-08-06 100.0 85.61 92.32 93.06 97.0 92.42 86.76 90.22 86.28 88.5 92.12
gpt-4-0125-preview 100.0 86.9 90.96 90.42 94.44 87.97 82.34 94.11 85.46 84.7 87.57
gpt-4-1106-preview 97.5 87.22 92.28 91.8 90.65 85.59 82.5 95.65 84.95 85.6 96.66
haiku 80.35 58.8 64.91 67.91 55.78 72.39 58.07 68.76 57.09 56.52 76.37
sonnet3 68.22 59.53 64.76 64.16 56.56 60.68 58.91 67.34 62.31 54.51 57.28
opus 100.0 74.56 81.16 84.86 83.96 83.13 73.58 88.64 80.1 79.47 92.12
sonnet35 95.0 76.45 76.7 82.36 82.74 80.64 77.01 85.68 77.95 80.98 88.79
mistralnemo 80.72 69.46 78.49 74.86 72.83 77.66 62.26 79.63 71.68 64.82 80.91
mistralsmall 73.22 64.58 73.46 72.36 70.05 77.91 56.36 71.89 66.32 62.75 75.16
mistral-large-2402 63.57 56.19 64.19 61.94 55.78 62.71 41.14 63.52 60.79 50.95 42.72
mistrallarge 95.0 82.87 88.18 93.2 86.53 88.9 79.13 88.68 82.45 84.77 92.12
llama3-8b 80.35 66.7 73.79 74.86 66.83 78.02 57.75 78.06 76.04 60.07 71.82
llama3-70b 97.5 78.54 87.9 87.64 85.74 85.37 75.94 86.34 85.66 78.65 79.7
llama3-1-8b 85.35 69.87 73.92 70.84 70.4 77.77 58.46 84.07 74.86 62.38 69.7
llama3-1-70b 97.5 76.29 88.43 87.78 86.53 83.02 76.98 91.65 87.39 80.15 92.12
llama3-1-405b 100.0 80.23 85.31 86.11 88.31 88.8 80.97 89.56 84.36 88.24 87.57
gemma2-9b 87.85 70.11 82.81 86.53 78.96 81.18 68.54 81.08 78.52 67.92 76.37
gemma2-27b 87.85 72.66 84.63 85.0 84.31 85.03 69.07 90.59 80.03 71.28 76.37
mistral-7b-v1 61.43 53.36 57.2 56.25 58.92 68.35 49.33 63.76 60.71 55.0 56.06
mistral-7b-v2 73.22 51.56 55.45 57.64 54.14 66.8 43.44 58.6 54.08 47.24 60.61
mixtral-8x22B 80.72 69.83 78.36 80.0 75.4 81.04 64.61 83.49 74.98 68.13 76.37
qwen1-5-72b-chat 97.5 74.54 85.48 90.56 82.18 83.03 71.01 88.01 81.32 76.98 80.91
qwen2-72b 97.5 81.45 86.94 86.39 89.87 85.78 69.86 87.59 78.14 83.03 95.46
random 24.28 29.85 22.76 26.39 32.39 28.26 25.32 20.8 25.59 22.9 25.76
human 91.65 92.27 93.6 92.92 92.31 91.65 92.53 94.48 93.47 92.08 90.06
@@
Table 5: Accuracy (%) for 30 tested models on CulturalBench-Easy at continent level.
Appendix F Evaluation results on CulturalBench-Hard
Models North America South America East Europe South Europe West Europe Africa Middle East/ West Asia South Asia Southeast Asia East Asia Oceania
gpt-3-5-turbo-1106 36.78 26.45 20.21 29.58 37.4 28.37 17.61 36.74 25.33 23.38 36.97
gpt35turbo 41.78 36.27 30.65 34.03 43.53 34.22 21.0 46.51 30.28 28.76 36.97
gpt4omini 63.57 45.54 37.04 50.0 55.56 48.76 37.87 60.13 50.08 47.44 44.84
gpt4o 70.72 54.19 54.3 65.69 68.25 67.76 54.22 72.5 64.49 61.38 54.84
gpt-4o-2024-08-06 75.72 55.76 52.73 64.03 69.81 65.2 59.05 68.42 61.39 62.49 60.61
gpt-4-0125-preview 75.72 57.42 56.57 60.14 68.03 67.56 54.94 66.7 59.99 65.22 59.39
gpt-4-1106-preview 77.85 52.82 53.9 61.25 60.12 66.87 53.23 73.5 59.56 61.55 56.06
haiku 43.93 30.17 28.36 20.84 33.96 27.64 19.42 33.0 17.2 20.32 27.88
sonnet3 48.93 31.01 33.1 52.08 40.09 34.68 26.3 40.47 26.67 32.63 30.3
opus 75.35 43.52 47.59 54.86 52.78 58.94 42.45 64.37 58.28 52.95 53.94
sonnet35 70.72 40.32 42.9 58.61 59.92 57.56 43.69 63.64 55.25 49.63 65.16
mistralnemo 48.93 30.09 37.84 46.94 45.66 39.86 29.76 45.78 39.63 34.06 34.84
mistralsmall 63.22 43.4 39.54 46.25 48.0 53.2 30.37 60.24 44.62 33.36 36.97
mistral-large-2402 63.22 39.53 50.15 52.22 42.3 52.84 40.46 62.88 58.72 44.79 48.48
mistrallarge 72.85 34.93 44.11 49.58 48.88 42.08 39.13 45.09 51.78 44.22 42.72
llama3-8b 24.28 20.56 15.56 26.25 14.48 19.67 10.57 26.67 25.43 26.05 19.09
llama3-70b 27.14 26.2 24.09 23.75 34.52 30.02 27.92 38.54 30.01 33.73 41.52
llama3-1-8b 51.43 33.81 37.35 44.58 34.3 38.75 28.74 47.43 40.85 37.02 40.61
llama3-1-70b 75.72 46.18 45.53 56.25 63.69 60.15 50.04 62.72 51.5 57.19 53.94
llama3-1-405b 73.22 50.27 47.05 57.92 54.34 55.01 42.71 64.46 48.91 54.15 49.39
gemma2-9b 51.43 37.71 30.69 40.84 40.08 42.89 29.62 54.04 39.62 36.24 16.66
gemma2-27b 65.72 42.25 37.01 38.2 55.35 50.65 32.82 60.94 36.51 40.82 30.3
mistral-7b-v1 31.78 26.2 18.53 22.08 22.39 22.72 11.43 21.71 15.67 16.68 15.76
mistral-7b-v2 41.78 28.49 26.45 34.03 36.96 38.46 27.84 47.3 38.88 33.22 33.64
mixtral-8x22B 68.22 37.08 39.08 44.3 53.01 50.73 37.73 66.94 43.24 38.46 38.18
qwen1-5-72b-chat 75.35 45.88 40.48 45.84 44.22 49.83 29.83 65.22 44.87 41.15 41.52
qwen2-72b 75.35 40.44 44.23 50.7 54.0 49.83 37.52 55.04 49.48 50.72 55.16
random 21.78 5.7 9.98 5.28 3.34 4.95 8.01 8.07 5.2 5.92 3.34
human 94.0 91.26 92.58 94.26 92.84 94.0 91.48 94.29 92.2 92.57 93.29
Table 6: Accuracy (%) for 30 tested models on CulturalBench-Hard at continent level.
Appendix G Annotators’ demographics in prolific platform
@@
North America South America East Europe South Europe West Europe Africa Middle East/ West Asia South Asia Southeast Asia East Asia Oceania
Gender (%)
Female 52.76 41.73 55.26 43.21 32.5 56.22 50 39.47 57.26 65.79 53.27
Male 46.62 58.26 44.74 56.79 67.5 42.49 48.95 59.87 42.74 32.89 46.73
Prefer not to say 0.61 0 0 0 0 1.29 1.05 0.66 0 1.32 0
Age (%)
$\leq$ 29 28.83 40.08 55.79 40.74 40.35 52.36 52.63 54.62 23.91 45.39 32.71
30-39 26.38 40.08 28.42 24.89 29.82 34.44 35.79 39.47 30.43 34.87 32.71
40-49 27.61 14.46 12.63 25.93 16.37 8.58 6.32 4.61 32.61 13.16 23.36
50-59 14.11 3.72 3.16 6.17 8.77 4.29 1.58 1.32 21.74 6.58 5.61
60-69 1.84 1.65 0 2.47 4.09 0.43 3.68 0 26.09 0 4.67
70-79 1.23 0 0 0 0.58 0 0 0 43.48 0 0.93
Student status (%)
No 85.28 64.88 56.84 51.85 76.02 49.79 45.26 48.03 60.48 61.18 66.36
Yes 9.82 30.99 37.89 46.91 22.22 45.92 45.26 46.71 31.05 23.68 24.30
Table 7: Annotators demographic in Prolific for the whole dataset before filtering. We set two main recruitment criteria to ensure the recruited annotators have a deep understanding of culture of the targeted country or region: (1) Nationality. (2) Primary residence before age 18. See the detail in Section 3.2.
Appendix H AI-assisted Red teaming System

### v2

@@ Acknowledgments
Appendix A Supplementary detail and experiments
Criteria on selecting underrepresented region/culture. Our goal in CulturalBench is to cover diverse regions/cultures that are typically neglected in LLM evaluation and training due to factors such as the dominance of English-based internet data and researcher geographical distribution. We consider "underrepresented" regions as those with significant populations but have disproportionately low representation in LLM development. In practice, beyond English-speaking regions (primarily the United States) and mainland China, many regions qualify as "underrepresented." While we didn’t apply strict criteria for defining "underrepresented" regions, we used heuristics including country/region populations and numbers of native language speakers to build our initial set of regions. We then filtered for regions with at least 10 active annotators on Prolific (our annotation platform vendor) as our validation protocol requires a minimum of 5 annotators from each region.
Additional Multilingual Experiment. We conducted an additional experiment with translated questions. We collaborated with 3 native volunteers to translate questions related to China (Simplified Chinese), Taiwan (Traditional Chinese), and Hong Kong (Cantonese). We then evaluated two models: GPT-4o and Qwen 2.5 72B (from a Chinese LLM provider). The results are summarized below:
Model Full Single-Mode Multi-Mode
GPT-4o (English subset) 90.6% 91.3% 80.0%
GPT-4o (Chinese) 87.9% 88.4% 80.0%
Qwen 2.5 72B (English subset) 85.9% 85.5% 90.0%
Qwen 2.5 72B (Chinese) 79.9% 79.7% 80.0%
Table 2: Easy Version (Accuracy %): Performance comparison across subsets and mode types.
Model Full Single-Mode Multi-Mode
GPT-4o (English subset) 62.4% 63.3% 50.0%
GPT-4o (Chinese) 57.0% 57.6% 50.0%
Qwen 2.5 72B (English subset) 55.7% 54.7% 70.0%
Qwen 2.5 72B (Chinese) 52.3% 52.5% 50.0%
Table 3: Hard Version (Accuracy %): Performance comparison across subsets and mode types.
Across both models, performance on translated questions is slightly lower than performance on the English versions, though the differences are not dramatic. This suggests that our English-only approach provides a reasonable assessment of cultural knowledge while avoiding potential confounds from language proficiency variations.
Another way of evaluation: Distribution-based Evaluation We explored using GPT-4o to calculate probability distributions over answers, computing linear probability using log probabilities. Our analysis of these linear probabilities revealed a highly right-skewed distribution: In total, there are 6784 instances (binary statement questions) for 1696 questions.
  • (50-60%]: 1.7% of responses
  • (60-70%]: 1.8% of responses
  • (70-80%]: 2.0% of responses
  • (80-90%]: 2.7% of responses
  • (90-100%]: 26.4% of responses
  • Exactly 100%: 65.4% of responses
This highly right-skewed distribution reveals that GPT-4o is typically extremely confident in its answers (>90% confidence in 91.78% of cases). This overconfidence suggests that using probability distributions as an evaluation metric may not meaningfully change the conclusions, as models rarely express uncertainty even when incorrect.
While we appreciate the reviewer’s suggestion about rephrasing questions to provide more context, we believe the current format effectively reveals important limitations in models’ cultural understanding - specifically their tendency to provide singular, overly confident answers in scenarios where cultural nuance requires acknowledging multiple valid perspectives.
Paper Task Will it be listed in which table for the revised version?
CultureBank: An Online Community-Driven Knowledge Base Towards Culturally Aware Language Technologies Shi et al. (2024) Knowledge base + automatic evaluation Not specifically for QA task - Extensive table
Auditing and mitigating cultural bias in llms//Cultural Bias and Cultural Alignment of Large Language Models Using WVS to evaluate on different cultural prompting variations Tao et al. (2023) Not QA eval - NA
DLAMA: A Framework for Curating Culturally Diverse Facts for Probing the Knowledge of Pretrained Language Models Keleg and Magdy (2023) Knowledge base + Probing Not specifically for QA task - Extensive table
Towards Measuring the Representation of Subjective Global Opinions in Language Models Durmus et al. (2023a) Use WVS/PEW to measure representation (percentage) of people’s opinions. Not specifically for QA task - Extensive table
Knowledge of cultural moral norms in large language models Ramezani and Xu (2023) Probe LMs to get cultural moral norms Not QA eval - NA
Towards measuring and modeling" culture" in llms: A survey Adilazuarda et al. (2024) Not QA eval - NA
CultureLLM: Incorporating Cultural Differences into Large Language Models Li et al. (2024) Prompting the LLMs to reduce cultural bias. Not QA eval - NA
Massively multi-cultural knowledge acquisition Fung et al. (2024) lm benchmarking & Building knowledge base of cultural knowledge + evaluation Not specifically for QA task - Extensive table
(Candle:) Extracting Cultural Commonsense Knowledge at Scale Nguyen et al. (2022) Building knowledge base + evaluation Not specifically for QA task - Extensive table
BLEND: A Benchmark for LLMs on Everyday Knowledge in Diverse Cultures and Languages Myung et al. (2024) Text-based short-form/long-form QA Continue to keep in Table 1
NormAd: A benchmark for measuring the cultural adaptability of large language models Rao et al. (2024) Text-based short-form QA Continue to keep in Table 1
FORK: A Bite-Sized Test Set for Probing Culinary Cultural Biases in Commonsense Reasoning Models Palta and Rudinger (2023b) Text-based short-form QA Newly added to Table 1
Eticor: Corpus for analyzing llms for 747 etiquettes. Dwivedi et al. (2023) Text-based subjective preference/judgement eval Not specific to QA - Extensive table
Having beer after prayer? measuring cultural bias in large language models. Naous et al. (2023) Analysis model response/ wvs Not QA eval - NA
CVQA: Culturally-diverse Multilingual Visual Question Answering Benchmark Romero et al. (2024) Visual QA Newly added into Table 1
CulturalVQA Benchmarking Vision Language Models for Cultural Understanding Nayak et al. (2024) Visual open-end QA from Candle Involving automatic eval and less relevant to text-based QA - Extensive table
CaLMQA: Exploring culturally specific long-form question answering across 23 languages Arora et al. (2024) Text-based long-form QA + automatic eval Automatic eval - Extensive table
BERTAQA: How Much Do Language Models Know About Local Culture?Etxaniz et al. (2024) Text-based short-form trivia QA Newly added into Table 1
CulturalBench (Our Work) Text-based short-form QA Table 1
Table 4: (Expanded table) Related Work on Cultural Awareness in Language Models
Extension table for related work. In Table 1, we focus specifically on QA benchmarks. Here, we incorporated additional relevant wor to provide a more comprehensive comparison in Table 4.
Appendix B Definition, examples and prompts for cultural topics
@@
Prompt for classification:
What is the most relevant topic for this question? Topics: <topic description and examples above> Instruction: Choose only one most relevant topic among the 17 topics. Only output the topic name e.g. Clothing, Food. Question: <Question>
@@
Table 5: Examples of CulturalBench for each topic. (Part 1)
@@
Table 6: Examples of CulturalBench for each topic. (Part 2)
Appendix C CulturalBench statistics
Country Counts
North Africa ($N=75$)
Morocco 38
Egypt 37
South Africa ($N=97$)
Zimbabwe 39
South Africa 58
West Africa ($N=34$)
Nigeria 34
North America ($N=54$)
United States 35
Canada 19
South America ($N=188$)
Chile 35
Mexico 49
Brazil 33
Peru 36
Argentina 35
East Asia ($N=243$)
South Korea 41
Japan 53
Taiwan 54
Hong Kong 36
China 59
South Asia ($N=156$)
India 46
Bangladesh 45
Nepal 33
Pakistan 32
Table 7: Distribution of questions across 45 countries in CulturalBench (part 1)
Country Counts
Southeast Asia ($N=203$)
Vietnam 33
Malaysia 35
Philippines 44
Indonesia 32
Singapore 32
Thailand 27
Middle East/West Asia ($N=171$)
Iran 37
Israel 42
Lebanon 37
Saudi Arabia 17
Turkey 38
East Europe ($N=138$)
Ukraine 34
Czech Republic 32
Romania 38
Poland 34
North Europe ($N=62$)
United Kingdom 33
Russia 29
South Europe ($N=76$)
Spain 40
Italy 36
West Europe ($N=130$)
Netherlands 51
France 47
Germany 32
Oceania ($N=69$)
New Zealand 32
Australia 37
Table 8: Distribution of questions across 45 countries in CulturalBench (part 2)
North Africa South Africa West Africa North America South America East Asia South Asia Southeast Asia Middle East/ West Asia East Europe North Europe South Europe West Europe Oceania
count (%) 5.17 5.47 2.34 9.28 11.43 5.32 9.06 12.49 9.55 9.28 2.49 3.06 7.77 7.28
Ethnicity(%)
Asian 0.0 2.07 0.0 19.92 0.33 98.58 91.67 94.86 7.91 0.41 60.61 4.94 7.28 21.76
Mixed 29.2 2.76 4.84 6.91 31.35 0.71 7.08 2.72 13.44 1.22 0.0 2.47 6.8 15.54
Other 27.01 0.0 0.0 7.32 22.77 0.0 0.83 2.11 33.99 0.81 3.03 2.47 0.49 5.18
White 42.34 10.34 0.0 42.28 44.88 0.0 0.42 0.3 44.27 97.56 34.85 88.89 83.01 57.51
Age group(%)
18-29 68.61 40.0 46.77 26.83 41.91 47.52 50.42 56.8 52.96 53.25 15.15 40.74 50.97 31.61
30-39 29.93 33.1 46.77 27.64 39.93 33.33 41.67 25.98 33.99 31.71 28.79 24.69 31.07 28.5
40-49 0.73 18.62 4.84 26.83 13.86 13.48 6.67 14.8 6.72 10.98 25.76 25.93 12.14 24.87
50-59 0.0 6.9 1.61 13.01 2.97 5.67 1.25 2.42 2.77 4.07 22.73 6.17 2.43 7.25
60 above 0.73 1.38 0.0 5.69 1.32 0.0 0.0 0.0 3.56 0.0 7.58 2.47 3.4 7.77
Sex(%)
Female 30.66 75.86 35.48 50.0 40.59 68.79 40.83 53.78 46.25 52.03 53.03 43.21 29.61 50.78
Male 67.15 24.14 64.52 49.59 59.41 29.79 58.75 45.32 52.96 47.56 46.97 56.79 69.9 48.19
Prefer not to say 2.19 0.0 0.0 0.41 0.0 1.42 0.42 0.91 0.79 0.41 0.0 0.0 0.49 1.04
Student Status(%)
No 35.77 53.79 53.23 79.67 62.05 58.16 44.58 60.42 44.27 56.91 84.85 51.85 66.5 63.73
Yes 52.55 41.38 41.94 13.01 31.02 29.79 45.83 29.91 45.85 31.71 9.09 46.91 27.18 21.76
Table 9: Annotators demographic in Prolific for the whole dataset before filtering. We set two main recruitment criteria to ensure the recruited annotators have a deep understanding of culture of the targeted country or region: (1) Nationality. (2) Primary residence before age 18. See details in Section 2.2.
@@ Appendix D Zero-shot evaluation prompts
Our evaluation is to ask the model in multiple choice setting. The zero-shot prompt is as follows to ensure the model only outputs one label (A, B, C, or D). We set the output token to be 1 for most of the models except for the reasoning models (e.g. o1, o1-mini, o3-mini and Deepseek r1) since they are not allowed.
@@
Our evaluation is to ask the model in binary setting (True/False). Our prompt is as follow to ensure the model only output one label (True/False). We set the output token to be 2 for most of the models except for the reasoning models (e.g. o1, o1-mini, o3-mini and deepseek r1) since they are not allowed.
@@
Model Overall North Africa South Africa West Africa North America South America East Asia South Asia Southeast Asia Middle East/ West Asia East Europe North Europe South Europe West Europe Oceania
claude-3.5-haiku 80.07 80.0 81.44 88.24 80.7 76.19 81.07 81.53 85.78 72.51 76.81 84.62 84.21 77.69 81.43
claude-3.5-sonnet 87.28 88.0 89.69 91.18 84.21 84.13 90.12 85.99 89.71 83.63 84.78 87.69 90.79 88.46 85.71
claude-3-opus 82.65 78.67 85.57 91.18 85.96 78.84 81.07 85.35 83.82 78.36 83.33 83.08 90.79 82.31 82.86
claude-3-sonnet 66.94 64.0 69.07 85.29 73.68 58.73 68.31 72.61 65.2 64.91 71.01 61.54 69.74 62.31 70.0
cohere-aya-32b 74.91 70.67 79.38 82.35 77.19 74.6 70.37 82.17 72.06 72.51 77.54 80.0 85.53 69.23 71.43
cohere-aya-8b 64.77 57.33 79.38 64.71 71.93 59.79 64.2 74.52 64.71 63.74 62.32 67.69 64.47 54.62 64.29
deepseek-v3 84.29 81.33 91.75 85.29 82.46 81.48 86.83 85.35 84.31 77.78 84.78 81.54 93.42 85.38 80.0
deepseek-r1 86.93 85.33 88.66 88.24 82.46 87.83 89.71 87.26 85.29 83.63 84.78 86.15 92.11 88.46 85.71
gemma-2-27b 78.25 76.0 87.63 88.24 80.7 74.6 75.72 84.08 79.41 70.18 79.71 83.08 81.58 76.15 75.71
gemma-2-9b 75.73 72.0 82.47 85.29 77.19 72.49 73.66 78.98 75.98 71.35 80.43 80.0 84.21 69.23 72.86
gpt-3.5-turbo-0125 70.11 68.0 83.51 73.53 73.68 67.2 67.49 75.8 68.14 60.82 67.39 78.46 77.63 70.77 70.0
gpt-4o 88.8 90.67 89.69 97.06 85.96 85.71 89.71 88.54 87.25 86.55 89.13 90.77 93.42 90.77 88.57
gpt-4o-mini 82.3 77.33 88.66 88.24 78.95 80.42 81.48 87.9 83.82 76.61 81.16 84.62 92.11 80.77 75.71
o1 89.62 90.67 89.69 85.29 85.96 90.48 90.53 88.54 89.71 87.72 89.86 93.85 92.11 92.31 82.86
o1-mini 81.48 82.67 83.51 82.35 84.21 77.25 82.3 82.17 83.33 74.85 83.33 78.46 86.84 85.38 78.57
o3-mini 83.41 81.33 82.47 85.29 85.96 83.07 84.77 86.62 82.35 77.19 83.33 86.15 90.79 85.38 77.14
grok2 86.87 91.89 87.67 90.91 96.3 82.0 85.1 89.62 88.08 82.11 92.77 84.62 90.54 84.51 92.31
llama-3-70b 81.36 80.0 84.54 88.24 80.7 79.89 82.72 82.8 84.8 75.44 81.88 81.54 88.16 76.92 75.71
llama-3.1-405b 85.17 81.33 89.69 91.18 85.96 78.84 88.89 86.62 87.75 81.29 83.33 90.77 86.84 82.31 84.29
llama-3.1-70b 83.24 77.33 85.57 91.18 84.21 80.95 82.72 89.17 85.29 78.36 84.78 83.08 85.53 80.77 81.43
llama-3.1-8b 70.22 61.33 83.51 76.47 77.19 66.67 67.08 82.8 71.08 61.4 74.64 73.85 73.68 60.0 67.14
llama-3.3-70b 82.36 76.0 84.54 85.29 75.44 78.31 85.19 87.9 83.82 77.19 86.23 83.08 85.53 80.0 80.0
mistral-7b 65.01 65.33 81.44 79.41 73.68 61.9 61.32 71.34 64.22 60.23 64.49 64.62 71.05 55.38 61.43
mistral-large 82.83 78.67 84.54 91.18 84.21 78.31 85.6 86.62 79.41 82.46 81.16 83.08 90.79 80.0 84.29
mistral-nemo 71.92 69.33 81.44 76.47 77.19 69.84 70.37 84.08 69.12 64.91 73.91 72.31 75.0 66.15 67.14
mistral-small 78.96 76.0 87.63 85.29 84.21 77.78 77.37 85.35 78.43 70.18 77.54 83.08 88.16 73.08 80.0
mixtral-8x22b 73.8 74.67 80.41 91.18 80.7 68.25 71.6 82.8 73.53 67.25 73.91 73.85 78.95 68.46 72.86
mixtral-8x7b 73.92 72.0 84.54 82.35 77.19 69.31 75.31 81.53 69.61 70.18 68.84 70.77 81.58 75.38 68.57
qwen-2.5-72b 83.18 78.67 85.57 91.18 82.46 84.66 85.6 86.62 80.39 74.85 84.78 86.15 89.47 82.31 78.57
@@
Table 10: Accuracy (%) for 29 tested models on CulturalBench-Easy at sub-continent level. Human baseline is 92.4% and the random baseline is 25%.
@@
Model Overall North Africa South Africa West Africa North America South America East Asia South Asia Southeast Asia Middle East/ West Asia East Europe North Europe South Europe West Europe Oceania
claude-3.5-haiku 50.12 49.33 55.67 70.59 64.91 42.33 54.73 56.05 48.04 44.44 37.68 55.38 57.89 47.69 48.57
claude-3.5-sonnet 53.46 46.67 57.73 70.59 64.91 49.21 58.44 57.96 52.45 42.69 42.75 58.46 65.79 53.85 52.86
claude-3-opus 52.87 54.67 61.86 55.88 59.65 41.8 56.79 62.42 55.39 44.44 50.0 66.15 51.32 44.62 50.0
claude-3-sonnet 47.25 44.0 54.64 52.94 56.14 42.33 43.21 52.23 47.55 47.95 45.26 42.42 51.32 45.38 51.43
cohere-aya-32b 31.18 25.33 35.05 38.24 54.39 34.39 28.4 35.67 30.39 18.71 36.96 29.23 26.32 30.77 30.0
cohere-aya-8b 28.66 25.33 29.9 35.29 40.35 32.28 29.63 33.12 30.73 25.73 26.28 26.15 13.16 23.85 28.57
deepseek-v3 38.86 32.0 49.48 41.18 50.88 40.21 44.44 37.58 38.73 23.39 34.06 47.69 47.37 34.62 38.57
deepseek-r1 57.33 46.67 56.7 61.76 57.89 56.08 61.32 59.87 53.43 51.46 61.59 60.0 71.05 54.62 55.71
gemma2-27b 42.56 29.33 54.64 61.76 61.4 40.21 46.09 51.59 35.29 35.67 34.06 60.0 39.47 40.77 34.29
gemma2-9b 41.44 25.33 52.58 64.71 59.65 37.57 44.44 53.5 38.73 33.33 34.06 50.77 39.47 35.38 37.14
gpt-3.5-turbo-0125 34.53 29.33 34.02 47.06 54.39 34.92 32.1 42.04 31.86 23.39 28.26 43.08 39.47 39.23 34.29
gpt-4o 60.49 54.67 68.04 73.53 71.93 57.14 64.61 68.79 59.31 49.12 52.17 66.15 69.74 56.15 57.14
gpt-4o-mini 46.78 41.33 53.61 50.0 61.4 42.33 51.03 55.41 46.08 34.5 38.41 55.38 48.68 47.69 44.29
o1 61.37 54.67 62.89 52.94 57.89 51.32 70.78 61.78 63.24 58.48 61.59 66.15 61.84 61.54 62.86
o1-mini 49.59 41.33 54.64 50.0 59.65 47.62 53.09 58.6 48.04 40.35 43.48 53.85 53.95 46.92 51.43
o3-mini 56.51 52.0 53.61 70.59 68.42 50.79 63.37 61.78 59.31 43.27 50.0 61.54 59.21 57.69 55.71
grok2 54.0 34.38 47.46 66.67 44.44 50.0 59.57 61.11 59.09 49.49 56.94 56.82 50.88 50.77 54.17
llama-3-70b 47.13 42.67 52.58 61.76 63.16 45.5 50.62 55.41 43.63 40.35 40.58 60.0 39.47 40.77 45.71
llama-3.1-405b 51.93 37.33 63.92 73.53 64.91 48.15 56.38 59.87 50.0 40.94 47.1 61.54 55.26 43.85 51.43
llama-3.1-70b 54.57 46.67 63.92 50.0 64.91 49.21 62.96 58.6 51.47 47.95 48.55 60.0 63.16 48.46 54.29
llama31-8b 35.99 26.67 48.45 41.18 59.65 35.45 37.45 42.04 34.31 26.32 31.16 43.08 42.11 26.15 32.86
llama-3.3-70b 54.4 48.0 69.07 58.82 66.67 50.26 62.55 57.32 47.06 48.54 46.38 61.54 61.84 48.46 52.86
mistral-7b 33.94 20.0 44.33 47.06 49.12 33.86 34.16 43.31 31.86 25.15 31.16 43.08 27.63 30.0 32.86
mistral-large 47.6 38.67 49.48 35.29 59.65 41.27 52.67 45.86 53.92 40.94 49.28 53.85 52.63 44.62 42.86
mistral-nemo 36.05 29.33 47.42 35.29 61.4 26.46 35.8 42.68 34.8 30.41 31.88 47.69 43.42 29.23 38.57
mistral-small-3 41.97 38.67 46.39 32.35 59.65 26.46 46.91 45.86 40.2 35.67 45.65 53.85 53.95 39.23 40.0
mixtral-8x22b 44.72 40.0 45.36 52.94 59.65 42.33 40.57 51.59 46.57 39.18 40.15 44.62 42.11 50.0 48.57
mixtral-8x7b 21.16 13.33 24.74 35.29 26.32 15.34 25.93 27.39 18.63 16.37 15.94 16.92 26.32 23.85 21.43
qwen25-72b 46.72 38.67 49.48 44.12 56.14 38.1 58.85 53.5 46.57 39.18 41.3 55.38 48.68 42.31 38.57
Table 11: Accuracy (%) for 29 tested models on CulturalBench-Hard at sub-continent level. Human baseline is 92.4% and the random baseline is $(\frac{1}{2})^{4}=6.25\%$.
Appendix E CulturalTeaming: AI-assisted Red teaming System